OpenStudy (anonymous):

MEDAL GIVEN TO CORRECT ANSWER. A 2 kg laptop sits on the floor near a 4 kg jar of pennies. If the force of gravity between them is 3.42 10-10 N, how far apart are they? A. 2.1 m B. 1.6 m C. 1.2 m D. 1.8 m

OpenStudy (anonymous):

try the equation....... F= k[(m1*m2)]/(r)2 try to find the r= radius....

OpenStudy (anonymous):

hellooo..... doubtss are welcome....

OpenStudy (tukitw):

By Newton's Law of Gravitation, the force between two objects with masses of M and m, at a distance r between their centers of mass, is given by \[F=G \frac{ Mm }{ r^{ 2 } }\], where G is the universal gravitational constant with a value of \[6.672 \times 10^{-11}N m^{2} kg^{-2}\] Just solve the equation for r and plug in the given values in their respective spot.

OpenStudy (theeric):

@harishk 's method is sound! It's Newton's Law of Gravitation! \(F=G\dfrac{m_1m_2}{r^2}\) You know everything from the problem except for the gravitational constant, \(G\). That \(G\) should be whatever your teacher uses, or maybe if there is a \(G\) defined earlier in the problems. Otherwise, it could be something in the book that the course uses. Or you can look for it on the internet.

OpenStudy (theeric):

Or @tukitw can post it, hehe :)

OpenStudy (tukitw):

lol

OpenStudy (theeric):

So the algebra and calculation is all that is left.

OpenStudy (anonymous):

i keep getting an answer thats not a choice. i think im plugging it in wrong. can you show me how to work it out? @theEric or @tukitw

OpenStudy (theeric):

Is it maybe a problem with things like including \(\times 10^{-11}\)? And getting those signs right?

OpenStudy (anonymous):

yeah, think so @theEric

OpenStudy (theeric):

Did you round to the tens place? That's what you'll have to do. Why don't you show how you solved for \(r\)? You can draw it if you don't want to type it, and there's also the equation button!

OpenStudy (anonymous):

im on my phone so i dont have those buttons, but im not sure where to plug in the numbers. like how do you know what goes where? @theEric

OpenStudy (theeric):

Oh, sorry! Ask about a variable, and I'll tell you what it means :) You can also type it out, like F=G*m_1* m_2 / r^2 which would be \(\left( F=G\dfrac{m_1m_2}{r^2}\right)\) I don't know if you can see that math, though.

OpenStudy (anonymous):

im curious as to where to put the numbers. what does the G, m1 & 2, & r mean? @theEric

OpenStudy (theeric):

G is the gravitational constant, which was given by @tukitw as \(6.672\times 10^{-11}\). m1 and m2 are the two masses in question. And r is the distance in between the two masses, and it is what you want to solve for.

OpenStudy (anonymous):

6.672 x 10 ^-11? @theEric

OpenStudy (theeric):

That is for G. It is just a number that makes the whole equation work.

OpenStudy (theeric):

It's called the gravitational constant, and it's important in the law of gravitation we're using.

OpenStudy (anonymous):

so if i dont have to solve for G, do i solve for m1 & 2, and r? @theEric

OpenStudy (theeric):

What does the question ask for? Feel free to copy and paste. You can also look for what is given to you in the problem.

OpenStudy (anonymous):

so, would i set it up 2-4/3.42? @theEric

OpenStudy (theeric):

Something like that, yes! Can you write out the whole equation?

OpenStudy (anonymous):

i thought that was. what am i missing? @theEric

OpenStudy (theeric):

it should be (2*4) and and the 3.42 is really 3.42*10^-11. Do you see why?

OpenStudy (anonymous):

i keep getting 2.3 and it's still not an answer. if i dont solve the 3.42 10 ^-11 then what am i doing wrong? @theEric

OpenStudy (theeric):

Have you tried to solve for \(r\) with algebra yet? That could be very helpful. You are missing G in your equation, also.

OpenStudy (anonymous):

i thought G was the 3.42 10 ^-11? @theEric

OpenStudy (theeric):

Nope! That's the force of gravitation - the force pulling the pennies and laptop together! I see how you could get confused there :) "gravitation" Do you want to try the algebra? After that, it should be very easy. F= G * m_1 * m_2 / r^2 You want to solve for r!

OpenStudy (anonymous):

so it's 3.42*2*4/? = the answer? @theEric

OpenStudy (theeric):

Close.. I think it might be a good idea to restart in your new question! :) Less messy, if you don't mind.

OpenStudy (anonymous):

ok folks lets clean this mess... \[F = G \frac{ m _{1} m _{2} }{ r ^{2} }\] here , m1 = mass 1 (let be laptop of 2 kg) m2 = mass 2 (let be pennies of 4 kg) F = force between two mass1 and mass 2 ( let be laptop and pennies as 3.42 * 10^-10 ) R is radius between two objects (which means distance .... you can take it as such....) and hell G is gravitational constant... so, the equation becomes..... \[r ^{2} = G \frac{ m _{1} m _{2} }{ F }\] so you will get it as \[r ^{2} = 6.672 * 10 ^{-11} \frac{ 2 * 4 }{ 3.42* 10 ^{-10} }\] so, finally you will get, \[r ^{2}\]= 1.5607 and r = root of 1.5607 = 1.2492 meters. Keep Asking & Answering.......